OpenStudy (anonymous):

Use a graph to solve the equation on the interval [−2π, 2π]. (List the solutions in increasing order from left to right on the x-axis.) sec x = sqrt 2

OpenStudy (anonymous):

Not with a graph but to help explain. \[\text{The secant of an angle is the multiplicative inverse of its cosine.}\\ \ \ \ \sec x = \frac1{\cos x}\\ \ \ \ \sec x = \sqrt 2\\ \ \ \ \cos x = \frac{\sqrt 2}2\\ \text{We know the unit circle identity states...}\\ \ \ \ \sin^2 x+\cos^2 x=1\\ \ \ \ \sin^2 x+(\frac{\sqrt 2}2)^2=1\\ \ \ \ \sin^2 x+\frac12=1\\ \ \ \ \sin^2 x=\frac12\\ \ \ \ |\sin x|=\frac{\sqrt 2}2\\ \text{Now we know the angles have a positive cosine which is equal}\\ \text{to the absolute value of the sine. What angles satisfy this equality? (think)} \ \ \ \frac\pi4,\frac{7\pi}4 \]

OpenStudy (anonymous):

Sorry I'm still confused. I understand a little bit with the unit circle, however, how do I make the equation: \[\left| \sin x \right| = \frac{ \sqrt{2} }{ 2 } \] into (x,y). Thanks!